OpenStudy (anonymous):

What is the weight (in grams) of a liquid that exactly fills a 182.8 milliliter container if the density of the liquid is 0.135 grams/milliliter? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

D = M/V 0.135 = M/182.8 now solve for M

OpenStudy (anonymous):

how am i supposed to do that

OpenStudy (anonymous):

nvm this question i have another

OpenStudy (anonymous):

M= 24.678 for future reference